OpenStudy (anonymous):

What is the radius of the circle given by the equation x^2+(y-1)^2=14 Round your answer to the nearest thousandth

OpenStudy (mathmate):

Hint: The standard equation of the circle centred on (x0,y0) with radius r is: (x-x0)^2 + (y-y0)^2 = r^2

OpenStudy (anonymous):

ok but how does that relate to the equation how do we get (x0,y0)

OpenStudy (mathmate):

x^2+(y-1)^2=14 can be written as (x-0)^2+(y-1)^2=14 so (x0,y0)=(0,1) and make your deductions for r^2.

OpenStudy (anonymous):

r^2=14?

OpenStudy (anonymous):

do you square root it?

OpenStudy (mathmate):

Exactly, to get the radius.

OpenStudy (anonymous):

\[2\sqrt{7}\]?

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Not really, you are "pushing" it! :) Sqrt(14) is already the simplest you can get, since the factors of 14 do not make perfect squares.

OpenStudy (anonymous):

oh yea..so is 14 my answer? How do i round my answer to the nearest thousandth?

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Yes sqrt(14) is your answer. To round to the nearest thousand, first find sqrt(14) using a calculator: sqrt(14)=3.741657.... A thousandth is the 3 place after the decimal. Draw a little line after the third place: 3.741 | 657 If the next digit to the right of the line is 4 or less, just drop everything to the right of the line. If the next digit to the right of the line is 5 or more, then drop everything to the right AND add one to the digit to the left of the line. Here the digit to the right is 6, which is greater than 5, so we drop the rest and add on to the digit on the left to get 3.742 as the answer.

OpenStudy (anonymous):

ok thank you!

OpenStudy (mathmate):

yw! :)